A proton described a circle of radius 1 cm in a magnetic field of strength 0.10 T. What would be the radius of the circle described by an α-particle moving with the same speed in the same magnetic field?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. r = mv/qB 0.01 = mv/e0.1 …(1) r = 4m * V/2e * 0.1 …(2) (2) ÷ (1) ⇒ r/0.01 = 4mVe * 0.1/2e * 0.1 * mv = 4/2 = 2 ⇒ r = 0.02 m = 2 cm.